Suppose that Alicia wants to buy a car. The dealer offers a financing package consisting of a 6% APR compounded monthly for a term of five years. Suppose that you want your monthly payments to be at most $320. What is the maximum amount that you should finance? Round your answer to the nearest dollar.
Answer by Theo(13342)

the present value of monthly payments of 320 at 6% apr for 5 years is equal to 16552.17944.

that's the most that she should finance if she doesn't want her payments to be greater than $320 per month.

i used a financial calculator to figure out the amount.

entries were:

number of time periods = 12 * 5 = 60
monthly interest rate = .06/12 = .005
monthly payment = $320.
solve for present value of a payment.

if you need the formulas, let me know and i'll get them for you.

you can also use an online calculator which i'll send you the link for if you need it.